7 times the sine of a certain angle is 7 times of the square of the cosine of the same angle. Find the angle in degrees if it’s in the first quadrant.
Answer #1

7 sino = 7 costo Sino costo 1-sino Sino= tu Sino + sind-1=0 (2) -(1) z using Dharacharya formula Sind=-b62-4ac -(2) аа Corepa7 sino = 7 costo Sino costo 1-sino Sino= tu Sino + sind-1=0 (2) -(1) z using Dharacharya formula Sind=-b62-4ac -(2) аа CorepaSino a OR = -1 + 2.23 2. Sino1-2-23 2 Sino OR 1:23 2 23 Sinop 3.23 2 Sino - 006) OR Sina - ): 61 0 = sin(0.61) OR Sin! (161)